The following table gives the available projects for a firm.

                                    A          B          C         D         E

Initial Investment         10        13        6          7          2

NPV                             3.5       –2        0.5       6          3

The firm has only $20 million to invest. Which project(s) will be accepted?
Answer #1

Accept Project D, Project E and Project A

Total investment = 7+2+10 = 19

Total NPV = 6+3+3.5 = 12.5